2 runners initially standing next to each other run away from each other with the speed,v and after 360m they are 360m apart. what is the velocity?
1
Expert's answer
2012-07-12T08:27:43-0400
The statement "after 360m they are 360m apart" is tautology, we'll appreciate it as a typo and understand as "after T seconds they are 360m apart". So,

2v·T = 360 ==> v = 360/(2T) = 180/T.